OpenStudy (anonymous):

solve for the value of "x" when cosA=3.5x and sinA=5.5x

OpenStudy (anonymous):

\[ \cos^2{A}=3.5^2x^2 \text{and } \sin^2{A}=5.5^2x^2 \\ \cos^2{A}+\sin^2{A}=1\\3.5^2x^2+5.5^2x^2=1 \]

OpenStudy (anonymous):

alternately:

OpenStudy (anonymous):

Hi @Algebraic! , meet again :D

OpenStudy (anonymous):

heya. what's for breakfast?

OpenStudy (anonymous):

Hahaha, it's afternoon here ;) You?

OpenStudy (anonymous):

nothing special. was hoping for some food porn cybur from you. what was for lunch?

OpenStudy (anonymous):

It was "nasi campur". Difficult to describe.

OpenStudy (anonymous):

nasi goreng I know... gotta google nasi campur

OpenStudy (anonymous):

looks awesome.

OpenStudy (anonymous):

Let's stop spamming at this question XD (It taste really hot)

OpenStudy (anonymous):

Rijsttafel for one...

OpenStudy (anonymous):

meh, he's offline and already got two solution methods... lol

OpenStudy (anonymous):

Lol :D

OpenStudy (anonymous):

I wish I had this when I was undergrad&high school: wander by, post a problem you don't want to do, wander back by in a few hours, *poof* done.

OpenStudy (anonymous):

actually, I don't, because I probably wouldn't have learned anything...

OpenStudy (anonymous):

So, you are a graduate student i guess?

OpenStudy (anonymous):

Balinese version looks pretty extravagant. yeah I'm starting grad school.

OpenStudy (anonymous):

Wow, i hope i could learn much from you then :) (Yeah, foods here sometimes difficult swallow because too hot).

OpenStudy (anonymous):

"Wow, i hope i could learn much from you then :)" I wouldn't count on it lol. I'd have to know much to be able to teach much;)

OpenStudy (anonymous):

(^ ^)b

OpenStudy (anonymous):

@jasp

OpenStudy (anonymous):

thanks for answering.. !! :))